A sportsman had increased his average speed from 250 to 300 m/min and he was finishing the distance one min faster. What is the distance he was running?

Answer by stanbon(75887)

A sportsman had increased his average speed from 250 to 300 m/min and he was finishing the distance one min faster. What is the distance he was running?
---
Slower speed DATA:
rate = 250 m/min ; time = x minutes ; distance = r*t = 250x meters
-----------------
Faster speed DATA:
rate = 300 m/min ; time = x - 1 minutes ; distance = 300x - 300 meters
------------------------
Equation:
dist = dist
250x = 300x - 300
----------------------------
50x = 300
x = 6 minutes
----
Ans: distance = 250x = 250*6 = 1500 meters
---------------
Cheers,
Stan H.
----------